OpenStudy (anonymous):

Rationalize the denominator. square root of 8/7

OpenStudy (anonymous):

* 6/7

OpenStudy (saifoo.khan):

\[\sqrt \frac 67?\]

OpenStudy (anonymous):

got it. i got another problem that involves rationalizing the denom

OpenStudy (anonymous):

\[\frac{ 3 }{ \sqrt{x}-4 }\]